Michael bought some tiger fish T $3 each and some goldfish at $4 each for his new aquarium. If Michael bought a total of 12 fish and spent a total of $40, how many goldfish did he buy
Answer by harrysgirl143(11)

1. t+g=12
2. 3t+4g=40
g= -t+12
3t-t+12=40
2t+12=40
2t=22
t=11

11 tigerfish
1 goldfish